The members of a cooking club are making cakes, which they will sell at a street fair for $13 apiece. It cost $21 for a booth at the fair, and the ingredients for each cake cost $6. At some point, the club members will sell enough cakes so that their sales cover their expenditures. How much will the sales and expenditures be? How many cakes will they have sold?
Answer by richwmiller(17219)

13c=sales
expenditures =21+6c
sales will cover expenditures
sales = expenditures
13c=21+6c
7c=21
c=3 sales will cover expenditures when they have sold 3 cakes
21+6*3=21+18=$39
13c*3=$39 sales will cover expenditures at $39